Soybean meal is 18% protein; cornmeal is 9% protein. How many pounds of each should be mixed together in order to get 360-lb mixture that is 15% protein.
How many pounds of the cornmeal should be in the mixture?
How many pounds of the soybean should be in the mixture?

Answer by mananth(16949)

percent percentage ---------------- quantity
Soyabean 18 ---------------- x lb
cornmeal 9 ---------------- 360 - x lb
Mixture 15.00% ---------------- 360

18 x + 9 ( 360 - x ) = 15 * 360

18 x + 3240 - 9 x = 5400
18 x - 9 x = 5400 - 3240
9 x = 2160
/ 9
x = 240 lb 18.00% Soyabean
120 lb 9.00% cornmeal